TheJag18:

I need all the work for this (Credit and Debt Payoff HC) A new homeowner is purchasing a living room set for $2,784 and must decide between two monthly installment financing offers. Offer 1: $225 down payment, remaining balanced financed at a 24.90% interest rate for 3 years. Offer 2: $450 down payment, remaining balanced financed at a 22.90% interest rate for 4 years. Part A: What is the total cost of offer 1? Explain which technology you used to solve and each step of your process. (3 points) Part B: What is the total cost of offer 2? Explain which technology you used to solve and each step of your process. (3 points) Part C: Which financing offer should the new homeowner choose? Explain your reasoning. (4 points)

Phantomdex:

Part A: AmountFinanced=PurchasePrice−DownPayment Amount\Financed = $2,784 - $225 Amount\Financed = $2,559 TotalCost=AmountFinanced×(1+ n/r)^nt Amount\,Financed = $2,559 r= 24.90% = 0.249 r= 24.90% = 0.249 r=12 (compounded monthly) t=3 years Total Cost = 2559 x (1 + 0.249/12)^12 x 3 TotalCost ≈ 2559×(1+0.02075)^36 TotalCost ≈ 2559×(1.02075)^36 TotalCost ≈ 2559×(1.9231) Total Cost ≈ 4922.49 So, the total cost of offer 1 is approximately $4,922.49. Part B: AmountFinanced=PurchasePrice−DownPayment Amount\Financed = $2,784 - $450 Amount\Financed = $2,334 TotalCost=AmountFinanced × (1+ n/r) ^nt Amount\,Financed = $2,334 r = 22.90%=0.229 n = 12 (compounded monthly) t = 4 years TotalCost = 2334×(1+ 12/0.229)^12×4 TotalCost≈2334×(1+0.01908)^48 TotalCost≈2334×(1.01908)^48 TotalCost≈2334×(1.71708) TotalCost≈4006.84 So, the total cost of offer 2 is approximately $4,006.84. Part C: Comparing the total costs: Offer 1: $4,922.49 Offer 2: $4,006.84 The new homeowner should choose offer 2 because it has the lower total cost. Despite having a longer repayment period, the lower interest rate and larger down payment result in a lower overall cost compared to offer 1.

Phantomdex:

My brain hurts.

TheJag18:

mine to trying to understand lmao

Phantomdex:

Don't just plug in the numbers smh, trust me

TheJag18:

well i need to put this in my essay question.

LookingForLove45:

as soon as I saw this, I got a headache keep up the good work👍